OpenStudy (anonymous):

I am working on Analyzing Graphs of Quadratic Functions, the problem I am stuck on is f(x)=x^2/2+4x+4. I think the area I am messing up on is the squaring. Someone please putting this in the most simplest form. Algebra is my worst enemy.

OpenStudy (anonymous):

what exactly are you supposed to do?

OpenStudy (anonymous):

put it in the form f(x)=a(x-h)^2+k. find the vertex and graph it.

OpenStudy (anonymous):

is it \[y=\frac{x^2}{2}+4x+4\]?

OpenStudy (anonymous):

that's the original function. but I don't know how to work it out to because an order pair

OpenStudy (anonymous):

if you want to write in vertex form first factor out the \(\frac{1}{2}\) from the first two terms and write \[y=\frac{1}{2}(x^2+2x)+4\]

OpenStudy (anonymous):

easiest way to find the vertex is to use \(-\frac{b}{2a}\) for the first coordinate

OpenStudy (anonymous):

in your example it is \[-\frac{4}{2\times \frac{1}{2}}=-4\]

OpenStudy (anonymous):

i wrote something stupid \[y=\frac{1}{2}(x^2+2x)+4\] if you factor out the \(\frac{1}{2}\) you actually get \[y=\frac{1}{2}(x^2+\color{red}8x)+4\]

OpenStudy (anonymous):

thanks I got it now.